# Yongjiang Wua, Yongtao Lia, Lihua Fenga, Jiuqiang Liub, Guihai Yuc, [^2] Corresponding author

### v2

@@ Acknowledgement
Yongtao Li was supported by the Postdoctoral Fellowship Program of CPSF (No. GZC20233196). Lihua Feng was supported by the NSFC (Nos. 12271527, 12471022 and 12071484). Guihai Yu was supported by the NSFC (Nos. 12461062 and 11861019) and Natural Science Foundation of Guizhou (Nos. [2020]1Z001 and [2021]5609). Yongtao Li would like to thank Prof. Biao Wu and Dr. Zixiang Xu for the inspiring discussions on extremal set problems over the years.
@@ Appendix B The case $k=4$ in the Huang–Peng theorem
We proceed by applying induction on $n\geq 7$. The base case $n=7$ is trivial. Suppose that $n\geq 8$ and the result holds for integers less than $n$. Since $\mathcal{F}$ and $\mathcal{G}$ are shifted, by Lemma 2.2, we have $|\mathcal{F}(\bar{n})|\geq 3$. Clearly, $\mathcal{F}(\bar{n})$ is $2$-intersecting. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(\bar{n})$ are cross-intersecting. By induction hypothesis, we get
@@
Hence, $i\in\{x_{0},x_{1},x_{2}\}$ and $j\notin\{x_{0},x_{1},x_{2}\}$. By symmetry, we may assume that $x_{0}=i$. This implies that $|\mathcal{G}(i,j)|\leq 2n-7$. By the maximality of $|\mathcal{G}|$, we have $|\mathcal{G}(\bar{i},\bar{j})|=\binom{n-4}{2}$. Observe that $\mathcal{G}(i,\bar{j}),\mathcal{G}(\bar{i},j)\subseteq\left\{A\in\binom{[n]\backslash\{i,j\}}{3}:A\cap\{x_{1},x_{2}\}\neq\emptyset\right\}$. Let =$\mathcal{R}=\left\{A\in\binom{[n]\backslash\{i,j\}}{3}:\{x_{1},x_{2}\}\subseteq A\right\}.$ By the maximality of $|\mathcal{G}|$ and $\mathcal{G}\nsubseteq\mathcal{G}_{2}(n,4)$, we may assume that $\mathcal{R}\subsetneq\mathcal{G}(i,\bar{j})$ and $\mathcal{R}\subsetneq\mathcal{G}(\bar{i},j)$. Note that $\mathcal{G}(i,\bar{j})\cap\mathcal{G}(\bar{i},j)\subseteq\mathcal{R}$. So $(\mathcal{G}(i,\bar{j})\backslash\mathcal{R})\cap(\mathcal{G}(i,\bar{j})\backslash\mathcal{R})=\emptyset$. Since $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$ are cross-intersecting and $n-2\geq 7$, by Lemma B.4, we obtain

### v3

@@ Acknowledgement
We would like to thank Russ Woodroofe for sharing the reference [3], and Andrey Kupavskii for sharing [37]. We would also like to thank the referees for the valuable suggestions which greatly improved the presentation of the manuscript. Yongtao Li was supported by the Postdoctoral Fellowship Program of CPSF (No. GZC20233196). Lihua Feng was supported by the NSFC (Nos. 12271527 and 12471022). Guihai Yu was supported by the NSFC (Nos. 12461062 and 11861019) and Natural Science Foundation of Guizhou (Nos. [2020]1Z001 and [2021]5609).
@@ Appendix B The case $k=4$ in the Huang–Peng theorem
We proceed by applying induction on $n\geq 7$. For the case $n=7$, the upper bound in Lemma B.1 is $\binom{7}{3}$. Since $\mathcal{F}$ and $\mathcal{G}$ are cross-intersecting, at most half of the sets of $\binom{[7]}{4}\cup\binom{[7]}{3}$ belong to $\mathcal{F}\cup\mathcal{G}$. Thus $|\mathcal{F}|+|\mathcal{G}|\leq\binom{7}{3}$. Suppose that $n\geq 8$ and the result holds for integers less than $n$. Since $\mathcal{F}$ and $\mathcal{G}$ are shifted, by Lemma 2.2, we have $|\mathcal{F}(\bar{n})|\geq 3$. Clearly, $\mathcal{F}(\bar{n})$ is $2$-intersecting. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(\bar{n})$ are cross-intersecting. By the induction hypothesis, we get
@@
Hence, $i\in\{x_{0},x_{1},x_{2}\}$ and $j\notin\{x_{0},x_{1},x_{2}\}$. By symmetry, we may assume that $x_{0}=i$. This implies that $|\mathcal{G}(i,j)|=|s_{i,j}(\mathcal{G})(i,j)|\leq\binom{n-2}{2}-\binom{n-4}{2}=2n-7$. By the maximality of $|\mathcal{G}|$, we have $|\mathcal{G}(\bar{i},\bar{j})|=\binom{n-4}{2}$. Observe that $\mathcal{G}(i,\bar{j}),\mathcal{G}(\bar{i},j)\subseteq\left\{A\in\binom{[n]\backslash\{i,j\}}{3}:A\cap\{x_{1},x_{2}\}\neq\emptyset\right\}$. Let =$\mathcal{R}=\left\{A\in\binom{[n]\backslash\{i,j\}}{3}:\{x_{1},x_{2}\}\subseteq A\right\}.$ By the maximality of $|\mathcal{G}|$ and $\mathcal{G}\nsubseteq\mathcal{G}_{2}(n,4)$, we may assume that $\mathcal{R}\subsetneq\mathcal{G}(i,\bar{j})$ and $\mathcal{R}\subsetneq\mathcal{G}(\bar{i},j)$. Note that $\mathcal{G}(i,\bar{j})\cap\mathcal{G}(\bar{i},j)\subseteq\mathcal{R}$. So $(\mathcal{G}(i,\bar{j})\backslash\mathcal{R})\cap(\mathcal{G}(i,\bar{j})\backslash\mathcal{R})=\emptyset$. Since $\mathcal{G}(i,\bar{j})$ and $\mathcal{G}(\bar{i},j)$ are cross-intersecting and $n-2\geq 7$, by Lemma B.4, we obtain